# Interpretable Multi-Source Data Fusion Through Latent Variable Gaussian Process

### v1

@@ Abstract
With the advent of artificial intelligence (AI) and machine learning (ML), various domains of science and engineering communites has leveraged data-driven surrogates to model complex systems from numerous sources of information (data). The proliferation has led to significant reduction in cost and time involved in development of superior systems designed to perform specific functionalities. A high proposition of such surrogates are built extensively fusing multiple sources of data, may it be published papers, patents, open repositories, or other resources. However, not much attention has been paid to the differences in quality and comprehensiveness of the known and unknown underlying physical parameters of the information sources that could have downstream implications during system optimization. Towards resolving this issue, a multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed. The individual data sources are tagged as a characteristic categorical variable that are mapped into a physically interpretable latent space, allowing the development of source-aware data fusion modeling. Additionally, a dissimilarity metric based on the latent variables of LVGP is introduced to study and understand the differences in the sources of data. The proposed approach is demonstrated on and analyzed through two mathematical (representative parabola problem, 2D Ackley function) and two materials science (design of FeCrAl and SmCoFe alloys) case studies. From the case studies, it is observed that compared to using single-source and source unaware ML models, the proposed multi-source data fusion framework can provide better predictions for sparse-data problems, interpretability regarding the sources, and enhanced modeling capabilities by taking advantage of the correlations and relationships among different sources.
@@
Keywords multi-source modeling, experimental data, probabilistic machine learning, data fusion, interpretable machine learning
@@ 1 Introduction
To tackle the aforementioned challenges, the need for a multi-source data fusion that considers source variabilities becomes evident. By integrating data from diverse sources, multi-source data fusion modeling with interpretable ML can provide a more comprehensive and robust understanding of material behaviors, leading to faster discovery of novel materials. Specifically for experimental data sources, the data-driven ML approach can consider the inherent complexities and uncertainties associated with each experimental data resource through extraction of meaningful patterns and insights from different experimental conditions. This factor effectively mitigates the data availability and quality challenges. To achieve this, a Gaussian Process (GP) based multi-source modeling through Latent Variable Gaussian Process (LVGP) [18] is proposed to fuse numerous information (data) sources together to obtain a single representative interpretable prediction model. LVGP is novel ML model that allows incorporation of qualitative (categorical) variables in addition to quantitative (numerical) variables into surrogate modeling and has been implemented in variety of fields and applications such as adaptive materials design [19, 20, 21], non-hierarchical interpretable multi-fidelity modeling [22], and first-ever mixed-variable (continious and categorical) metamodel-based global sensitivity analysis [23]. Therefore, throughout this research, a single multi-source ML model that incorporates engineering system parameters as numerical (and categorical if there are any) variables while incorporating each data source as a qualitative variable is developed and studied. The qualitative information sources are incorporated into the LVGP model by implicitly mapping each qualitative variable into a low-dimensional quantitative latent space through statistical inference techniques. Specifically, as the influence of any qualitative variable on any quantitative response must be always due to some underlying (and possibly high-dimensional) quantitative physical variables, the latent variable approach provides physically meaningful dimension reduction. Here, the term “physically-meaningful” can be associated with explaining the cause-effect relationships between the information sources and the system response. Therefore, the qualitative representation of information sources indirectly accounts for the variability they bring through known and unknown underlying processing variables, which in return provides further interpretability regarding the information sources and their relationships on the system of interest.
@@ 2.1 Review of Gaussian Process Regression
With its powerful and versatile capabilities of capturing non-linear responses and providing fast and accurate predictive modeling with uncertainty quantification, Gaussian processes (GPs) are well-known metamodels that have been implemented in many engineering applications [24]. GPs are non-parametric regression models and are based upon Bayesian statistics in combination with numerous kernel methods to provide data-driven probabilistic predictions. Overall, GPs provide an immensely powerful tool for engineers seeking to build fast and accurate predictive models while quantifying and managing uncertainty.
@@
Although GPs are powerful ML models for numerous engineering applications, they tend to perform poorly when qualitative (categorical) variables are present in the input design space. The main challenge arises from the absence of a well-defined distance metric that differentiates qualitative variables. To overcome this vital challenge, Latent Variable Gaussian Process (LVGP) has been developed [18]. The details of the LVGP model are given in the next section.
@@ 2.2 Review of Latent Variable Gaussian Process
Due to the nature of the GP correlation functions, it is not plausible to directly include qualitative variables into the modelling since the difference between qualitative variables become unclear without a well-defined distance metric. Recently, Latent Variable Gaussian Process (LVGP) has been developed to integrate qualitative variables into GP modeling through quantitative latent variable representation of these variables to address this difficulty [18].
Consider a new input space, $\mathbf{b}=[\mathbf{x^{T}},\mathbf{t^{T}}]$ where both quantitative, $\mathbf{x}=[x_{1},x_{2},..x_{m}]\epsilon R^{m}$, and qualitative $\mathbf{t}=[t_{1},t_{2},...,t_{q}]\epsilon R^{q}$ input variables are included in this new domain. Each qualitative variable can have $j$ number of options (levels), i.e $t_{i}=[l_{1}(t_{i}),l_{2}(t_{i}),...,l_{j}(t_{i})]$ where $i=1,2,...,q$. For a given response of interest, $y(b)$, the influence of each qualitative variable can be represented by an underlying quantitative (numerical) space, $v_{t_{i}}(l_{j})=[v_{1}(l_{1}),v_{2}(l_{1}),...,v_{k}(l_{j})]\epsilon R^{k}$. This underlying quantitative space can possibly be undiscovered, unknown, complex, or high-dimensional. As a result, the main idea of LVGP lies in learning a low-dimensional quantitative latent space to approximate the original underlying space through statistical inference using the available input-output data. According to the original LVGP paper, a two-dimensional latent variable vector is sufficient to express the influence of the qualitative variables on the response of interest [18]. Therefore, qualitative variable, $t_{i}$ can be represented as $\mathbf{z}(t_{i})=[z_{1}(l_{1}),z_{2}(l_{1}),z_{1}(l_{2}),z_{2}(l_{2}),...,z_{1}(l_{j}),z_{2}(l_{j})]$, where each level $l_{j}(t_{i})$ is represented with a 2D latent vector $[z_{1}(l_{j}),z_{2}(l_{j})]$ Thus, the previously defined mixed-variable input space, $\mathbf{b}=[\mathbf{x}^{T},\mathbf{t}^{T}]$, then becomes $\mathbf{w}=[\mathbf{x}^{T},\mathbf{z(t)}{{}^{T}}]\epsilon R^{m+2\times q}$. With the newly defined input space, the previously specified GP correlation function (Equation 2) is updates as,
@@
With the introduction of LVGP, the primary contribution of this paper lies in the modeling of the multiple sources of information as a representative qualitative (categorical) input in addition to known quantitative (numerical) inputs through the LVGP model, making the surrogate modeling source-aware and interpretable. Further details of the multi-source data fusion modeling is provided in the subsequent section.
@@ 2.3 Interpretable Multi-Source Data Fusion through Latent Variable Gaussian Process
For an engineering system of interest, there could be multiple sources of information (patents, papers, open-source repositories) that aim to analyze the system. However, due to the diverse nature of data collection, each information source could contain different underlying procedures and processes that are not reported such as experimental errors, confidential information and so on. In this work, a multi-source modeling technique with Latent Variable Gaussian Process is proposed to combine all of the information sources into a single source-aware model that considers not only the relationships between information sources but also that both the known and unknown underlying physical parameters associated with each information sources when making predictions. To achieve this, a given engineering system of interest is modeled with a new input space, $[\mathbf{x},\mathbf{t},\mathbf{s}]$, where $\mathbf{x}$ and $\mathbf{t}$ are the known common numerical ($\mathbf{x}$) and categorical ($\mathbf{t}$) material parameters between information sources, and $\mathbf{s}$ is the newly introduced seperate categorical variable that represent the information sources ($\mathbf{s}$), along with the output response of interest, $\mathbf{y}$. Here, $\mathbf{s}$ is treated as a unique seperate qualitative variable that differentiates the information sources and integrates them into the multi-source model through the same statistical inference techniques used to estimate the latent variables of the variable $\mathbf{t}$. For this newly introduced variable $\mathbf{s}$, each information source is labelled with the letter $l$. A visual representation of the latent mapping of qualitative variable ($s$), which is denoted as the information sources where each source is represented by a level ranging from $l_{1}$ to $l_{j}$ in current case, is shown in Figure 1. By incorporating information sources as a qualitative variable, the aim is to learn and extract the unknown physical parametes underlying each information source through 2D latent spaces using statisitical inference techniques that has been reviewed in Section 2.2.1. It is important here to note that, the spatial relations (distances) between the information sources in the latent space can show similarity and differences regarding the influence of the information sources on the system of interest. Similarly, spatial relationships between latent variables can also imply the dimensionality of underlying known and unknown physical parameters related to the information sources. Therefore, for a given information source $j$, it is incorporated into the LVGP model with it’s unique known common parameters $x_{j}$ and $t_{j}$ and it’s unique source identifier $l_{j}$. Once the LVGP is trained, the model can become aware of not only the the known and unknown assumptions and parameters underlying each information source but also the relationships between each information sources. The awareness of the sources provided by the LVGP can further lead to fast and accurate predictions. Figure 2 shows the proposed multi-source modeling framework along with information source examples and their unique inputs to the LVGP model. It is important here to note that although common qualitative variables ($\mathbf{t}$) can be integrated into the multi-source model with LVGP, case studies with common quantitative variables ($\mathbf{x}$) are only demonstrated, leading to an input space with only $\mathbf{x}$ and $\mathbf{s}$ that are fed into the LVGP model.
@@
As previously mentioned, the latent variable approach provides physics-based dimension reduction that could potentially explain the "cause-effect" relationship between the inputs (the information sources) and outputs (response). Therefore, the proposed model can give us physical insights regarding the information sources by analyzing the spatial locations of the variables (sources) in the latent space. Here, the locations of the latent variables could suggest that the variables that are closer in the latent space can have similar impact on the response. Moreover, for the scenarios where no prior knowledge exists regarding information sources, the LVGP model can help us extract and reveal the relationships between the sources. To understand the underlying similarities and differences between information sources, a dissimilarity metric ($D$) that quantifies the dissimilarities between the sources through a distance metric is introduced. This dissimilarity metric is measured through the Euclidean distance of the data source $(z^{j})$ from the reference data source $(z^{*})$ in the latent space. The metric is given in Equation 7. For a given LVGP model, the latent variables are constrained to be between $z=[-3,3]$ as this was the typical practice in previous applications of LVGP and the reference data source will be placed at $z^{*}=[0,0]$ in the latent space. Therefore, the most dissimilar datasource from the reference datasource is going to be placed at [±$\pm 3,\pm 3$], making $z^{max}=[\pm 3,\pm 3]$. Based on this, the maximum distance between the reference and given information source can be $||z^{max}-z^{*}||=3\sqrt{2}$. Thus, a normalized unit of dissimilarity ($D$) can be calculated through Equation 8. With this quantified metric, the similarities and differences between information sources can be understood, which could be further used in downstream tasks such as anamoly detection and fault identification of sources.
Figure 2: Schematic of multi-source modeling through LVGP
@@ 3.1 Parabola Problem
Table 1: Performance for the parabola problem
@@
Figure 3: Summary of the Parabola problem. (a) Overview of Parabola Example (b-c) The parity plot between the GP and LVGP models (d-e) Prediction of Parabola on the Ground source GP model and LVGP model (f) Latent Space of Parabola Example from LVGP. Ground Source is denoted as the reference source (*) for dissimilarity metric, $D$, calculations
@@
The proposed method is next demonstrated on a well-known mathematical function that is used in optimization studies. Ackley’s function is a customiziable multi-dimensional function that has been implemented in numerous optimization applications for testing and validation purposes [25]. Here, 2D Ackley’s function, Equation 10, is implemented to further demonstrate the proposed interpretable multi-source data fusion method. In addition to the original 2D function, three (3) other variations of Ackley’s function are generated from the original function as different sources of information. Similar to the parabola problem, the original function is named as the Ground Source and the remaining 3 functions are named as Perturbed Sources 1, 2, and 3. The formulations of the functions are found in Supplementary Table 2 and their surface of the response, $z$, is shown in Figure 4a.
@@
Identical to the previous case study, two ML models are built (GP and LVGP) to demonstrate the effectivenes of the proposed approach. First, a GP model is built by only using the quantitative variables, $x$ and $y$ to predict the response, $z$. Then, in addition to the two quantative variables, a qualitative variable is added to differentiate the information sources for the LVGP model. This approach is the exact same approach taken for the previous parabola problem and will be implemented for the remainder of this work. During the training, 5-fold cross validation (CV) technique is implemeneted and tested on a separate testing set. Both the training and testing sets are created by sampling uniformly from the space $[x,y]\epsilon[-5,5]$. The number of training and testing samples employed to build and validate the models are avaiable in Supplementary Table 2. Next, a comparision study between the predictive performance of both models using the NMRSE metric is carried out. Table 2 (Model Types: GP and LVGP) shows the performance metrics for both models. From the results, it can be clearly observed once again that incorporating different sources as a seperate variable significantly improves the predicitive performance of the model. The model was also able to capture the relationships between the response surfaces and provide interpretability through the latent variables clearly. The latent space obtained from the LVGP, Figure 4b, match very well with the response surfaces, Figure 4a. Furthermore, the dissimilarity metric calculated using the Ground Source as the reference source, $D$, agree with the response surfaces and what is known regarding the similarity between the functions, demonstratng the efficancy of the method and the metric (Figure 4b).
@@
A possible follow up question one might ask as a response to the demonstrated results could be to build separate GP models to learn the information sources and predict their responses individually to achieve better accuracies. To answer this question, a sepearate GP model is built only using the Ground Source and compared with the previously built LVGP on a seperate testing set. For ease of notation, this single-source model is denoted as GP-GS. The trained GP-GS model only contained 20 samples from the Ground Source whereas the LVGP model contained all 170 samples from the available sources. The same testing set from the Ground Source is used and the prediction performances are compared. Table 2 (Model Types: LVGP and GP-GS) shows the prediction performances between the GP-GS and LVGP models, along with the parity plot in Figure 4c. In addition to achieving lower NRMSE values for accurate modeling, the parity plot demonstrates that the prediction uncertainity provided by the LVGP model is much less compared to the GP model. This result reveal an interesting phenomenon. Compared to the model that is only trained on the original single-source of information, gathering further information from other sources and explicitly denoting information sources into the data fusion modeling results in a more confident predictive model that provides lower predicition uncertainity on the original single-source of information.
@@
Table 2: Predictive performance comparision for GP, LVGP, and GP-GS models on the 2D Ackley function
Model   Type
Mean Training NRMSE
Mean CV NRMSE
Testing NRMSE
GP
$0.253$
$0.289$
$0.254$
LVGP
$0.013$
$0.029$
$0.024$
GP-GS
$0.089$
@@
The conclusions from this study can be significant for many reasons and can be applied to many engineering and scientific applications, specifically for materials science. This is because many researchers only have access or control over their own experiments/simulations and rely on building a surrogate model using their own data to guide their next property evaluations. On the other hand, due to the high costs of property evaluations, data curation for better predictive surrogate models could be expensive and limited to only a few samples. However, by incorporating more data from different available sources (open-sources, papers, patents) and explicitly differentiating each data source as an input to their models, researchers have the potential to obtain better and confident predictive model, which in return has the potential to lead to novel material candidate discovery at a faster rate. In the following two sections, Section 4 and Section 5, the advantages of using multi-source data fusion through LVGP on two materials design cases are demonstrated.
@@
Figure 4: (a) Response Surface of the Ground Source and Perturbed Sources for the 2D Ackley’s function. (b) The latent variables obtained from the LVGP model. Ground Source is denoted as the reference source (*) for dissimilarity metric, $D$, calculations (c) The parity plot between the GP-GS and LVGP models tested on the Ground Source data only

### v3

# references.bib

@@ Abstract
With the advent of artificial intelligence (AI) and machine learning (ML), various domains of science and engineering communities have leveraged data-driven surrogates to model complex systems from numerous sources of information (data). The proliferation has led to significant reduction in cost and time involved in development of superior systems designed to perform specific functionalities. A high proposition of such surrogates is built extensively fusing multiple sources of data, may it be published papers, patents, open repositories, or other resources. However, not much attention has been paid to the differences in quality and comprehensiveness of the known and unknown underlying physical parameters of the information sources that could have downstream implications during system optimization. Towards resolving this issue, a multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed. The individual data sources are tagged as a characteristic categorical variable that are mapped into a physically interpretable latent space, allowing the development of source-aware data fusion modeling. Additionally, a dissimilarity metric based on the latent variables of LVGP is introduced to study and understand the differences in the sources of data. The proposed approach is demonstrated on and analyzed through two mathematical (representative parabola problem, 2D Ackley function) and two materials science (design of FeCrAl and SmCoFe alloys) case studies. From the case studies, it is observed that compared to using single-source and source unaware ML models, the proposed multi-source data fusion framework can provide better predictions for sparse-data problems, interpretability regarding the sources, and enhanced modeling capabilities by taking advantage of the correlations and relationships among different sources.
@@
Keywords Latent Variable Gaussian Process, Gaussian Process Regression, Multi-Source Modeling, Data Fusion, Interpretable Artificial Intelligence, Uncertainity Quantification, Probabilistic Machine Learning
@@ 1 Introduction
To tackle the aforementioned challenges, the need for a multi-source data fusion that considers source variabilities becomes evident. By integrating data from diverse sources, multi-source data fusion modeling with interpretable ML can provide a more comprehensive and robust understanding of material behaviors, leading to faster discovery of novel materials. Specifically for experimental data sources, the data-driven ML approach can consider the inherent complexities and uncertainties associated with each experimental data resource through extraction of meaningful patterns and insights from different experimental conditions. This factor effectively mitigates the data availability and quality challenges. To achieve this, a Gaussian Process (GP) based multi-source modeling through Latent Variable Gaussian Process (LVGP) [zhang2020latent] is proposed to fuse numerous information (data) sources together to obtain a single representative interpretable prediction model. LVGP is novel ML model that allows incorporation of qualitative (categorical) variables in addition to quantitative (continuous) variables into surrogate modeling and has been implemented in variety of fields and applications such as adaptive materials design [iyer2020data, wang2020featureless, zhang2020bayesian], non-hierarchical interpretable multi-fidelity modeling [CHEN2024116773], and first-ever mixed-variable (continuous and categorical) metamodel-based global sensitivity analysis [comlek2023mixed]. Therefore, throughout this research, a single multi-source ML model that incorporates engineering system parameters as numerical (and categorical if there are any) variables while incorporating each data source as a qualitative variable is developed and studied. The qualitative information sources are incorporated into the LVGP model by implicitly mapping each qualitative variable into a low-dimensional quantitative latent space through statistical inference techniques. Specifically, as the influence of any qualitative variable on any quantitative response must be always due to some underlying (and possibly high-dimensional) quantitative physical variables, the latent variable approach provides physically meaningful dimension reduction. Here, the term “physically-meaningful” can be associated with explaining the cause-effect relationships between the information sources and the system response. Therefore, the qualitative representation of information sources indirectly accounts for the variability they bring through known and unknown underlying processing variables, which in return provides further interpretability regarding the information sources and their relationships on the system of interest.
@@ 2.1 Review of Gaussian Process Regression
With its powerful and versatile capabilities of capturing nonlinear responses and providing fast and accurate predictive modeling with uncertainty quantification, Gaussian processes (GPs) are well-known metamodels that have been implemented in many engineering applications [williams2006gaussian]. GPs are nonparametric regression models and are based upon Bayesian statistics in combination with numerous kernel methods to provide data-driven probabilistic predictions. Overall, GPs provide an immensely powerful tool for engineers seeking to build fast and accurate predictive models while quantifying and managing uncertainty.
@@
Although GPs are powerful ML models for numerous engineering applications, they tend to perform poorly when qualitative (categorical) variables are present in the input design space. The main challenge arises from the absence of a well-defined distance metric that differentiates qualitative variables. To overcome this vital challenge, Latent Variable Gaussian Process (LVGP) has been developed [zhang2020latent]. The details of the LVGP model are given in the next section.
@@ 2.2 Review of Latent Variable Gaussian Process
Due to the nature of the GP correlation functions, it is not plausible to directly include qualitative variables into the modelling since the differences between qualitative variables become unclear without a well-defined distance metric. Recently, Latent Variable Gaussian Process (LVGP) has been developed to integrate qualitative variables into GP modeling through quantitative latent variable representation of these variables to address this difficulty [zhang2020latent].
Consider a new input space, $\mathbf{b}=[\mathbf{x^{T}},\mathbf{t^{T}}]$ where both quantitative, $\mathbf{x}=[x_{1},x_{2},..x_{m}]\epsilon R^{m}$, and qualitative $\mathbf{t}=[t_{1},t_{2},...,t_{q}]\epsilon R^{q}$ input variables are included in this new domain. Each qualitative variable can have $j$ number of options (levels), i.e $t_{i}=[l_{1}(t_{i}),l_{2}(t_{i}),...,l_{j}(t_{i})]$ where $i=1,2,...,q$. For a given response of interest, $y(b)$, the influence of each qualitative variable can be represented by an underlying quantitative (numerical) space, $v_{t_{i}}(l_{j})=[v_{1}(l_{1}),v_{2}(l_{1}),...,v_{k}(l_{j})]\epsilon R^{k}$. This underlying quantitative space can possibly be undiscovered, unknown, complex, or high-dimensional. As a result, the main idea of LVGP lies in learning a low-dimensional quantitative latent space to approximate the original underlying space through statistical inference using the available input-output data. According to the original LVGP paper, a two-dimensional latent variable vector is sufficient to express the influence of the qualitative variables on the response of interest [zhang2020latent]. Therefore, qualitative variable, $t_{i}$ can be represented as $\mathbf{z}(t_{i})=[z_{1}(l_{1}),z_{2}(l_{1}),z_{1}(l_{2}),z_{2}(l_{2}),...,z_{1}(l_{j}),z_{2}(l_{j})]$, where each level $l_{j}(t_{i})$ is represented with a 2D latent vector $[z_{1}(l_{j}),z_{2}(l_{j})]$ Thus, the previously defined mixed-variable input space, $\mathbf{b}=[\mathbf{x}^{T},\mathbf{t}^{T}]$, then becomes $\mathbf{w}=[\mathbf{x}^{T},\mathbf{z(t)}{{}^{T}}]\epsilon R^{m+2\times q}$. With the newly defined input space, the previously specified GP correlation function (Equation 2) is updates as,
@@
With the introduction of LVGP, the primary contribution of this paper lies in the modeling of the multiple sources of information as a representative qualitative (categorical) input in addition to known quantitative (numerical) inputs through the LVGP model, making the surrogate modeling source-aware and interpretable. Further details of the multi-source data fusion modeling are provided in the subsequent section.
@@ 2.3 Interpretable Multi-Source Data Fusion Through Latent Variable Gaussian Process
For an engineering system of interest, there could be multiple sources of information (patents, papers, open-source repositories) that aim to analyze the system. However, due to the diverse nature of data collection, each information source could contain different underlying procedures and processes that are not reported such as experimental errors, confidential information and so on. In this work, a multi-source modeling technique with Latent Variable Gaussian Process is proposed to combine all of the information sources into a single source-aware model that considers not only the relationships between information sources but also that both the known and unknown underlying physical parameters associated with each information sources when making predictions. To achieve this, a given engineering system of interest is modeled with a new input space, $[\mathbf{x},\mathbf{t},\mathbf{s}]$, where $\mathbf{x}$ and $\mathbf{t}$ are the known common numerical ($\mathbf{x}$) and categorical ($\mathbf{t}$) system parameters between information sources, and $\mathbf{s}$ is the newly introduced separate categorical variable that represent the information sources ($\mathbf{s}$), along with the output response of interest, $\mathbf{y}$. Here, $\mathbf{s}$ is treated as a unique separate qualitative variable that differentiates the information sources and integrates them into the multi-source model through the same statistical inference techniques used to estimate the latent variables of the variable $\mathbf{t}$. For this newly introduced variable $\mathbf{s}$, each information source is labelled with the letter $l$. A visual representation of the latent mapping of qualitative variable ($s$), which is denoted as the information sources where each source is represented by a level ranging from $l_{1}$ to $l_{j}$ in current case, is shown in Figure 1. By incorporating information sources as a qualitative variable, the aim is to learn and extract the unknown physical parameters underlying each information source through 2D latent spaces using statistical inference techniques that has been reviewed in Section 2.2.1. It is important here to note that the spatial relations (distances) between the information sources in the latent space can show similarity and differences regarding the influence of the information sources on the system of interest. Similarly, spatial relationships between latent variables can also imply the dimensionality of underlying known and unknown physical parameters related to the information sources. Therefore, for a given information source $j$, it is incorporated into the LVGP model with it’s unique known common parameters $x_{j}$ and $t_{j}$ and it’s unique source identifier $l_{j}$. Once the LVGP is trained, the model can become aware of not only the known and unknown assumptions and parameters underlying each information source but also the relationships between information sources. The awareness of the sources provided by the LVGP can further lead to fast and accurate predictions. Figure 2 shows the proposed multi-source modeling framework along with information source examples and their unique inputs to the LVGP model. It is important here to note that although common qualitative variables ($\mathbf{t}$) can be integrated into the multi-source model with LVGP, case studies with common quantitative variables ($\mathbf{x}$) are only demonstrated, leading to an input space with only $\mathbf{x}$ and $\mathbf{s}$ that are fed into the LVGP model.
@@
As previously mentioned, the latent variable approach provides physics-based dimension reduction that could potentially explain the "cause-effect" relationship between the inputs (the information sources) and outputs (response). Therefore, the proposed model can give us physical insights regarding the information sources by analyzing the spatial locations of the variables (sources) in the latent space. Here, the locations of the latent variables could suggest that the variables that are closer in the latent space can have a similar impact on the response. Moreover, for the scenarios where no prior knowledge exists regarding information sources, the LVGP model can help us extract and reveal the relationships between the sources. To understand the underlying similarities and differences between information sources, a dissimilarity metric ($D$) that quantifies the dissimilarities between the sources through a distance metric is introduced. This dissimilarity metric is measured through the Euclidean distance of the data source $(z^{j})$ from the reference data source $(z^{*})$ in the latent space. The metric is given in Equation 7. For a given LVGP model, the latent variables are constrained to be between $z=[-3,3]$ as this was the typical practice in previous applications of LVGP and the reference data source will be placed at $z^{*}=[0,0]$ in the latent space. Therefore, the most dissimilar information source from the reference source is going to be placed at [±$\pm 3,\pm 3$], making $z^{max}=[\pm 3,\pm 3]$. Based on this, the maximum distance between the reference and given information source can be $||z^{max}-z^{*}||=3\sqrt{2}$. Thus, a normalized unit of dissimilarity ($D$) can be calculated through Equation 8. With this quantified metric, the similarities and differences between information sources can be understood, which could be further used in downstream tasks such as anomaly detection and fault identification of sources.
Figure 2: Schematic of multi-source data fusion through LVGP
@@ 3.1 Parabola Problem
Table 1: Predictive performance comparison between GP and LVGP for the parabola problem
@@
Figure 3: (a) Overview of Parabola Example. (b-c) The parity plot between the GP and LVGP models. (d-e) Prediction of Parabola on the Ground source GP model and LVGP model. (f) Latent Space of Parabola Example from LVGP. Ground Source is denoted as the reference source (*) for dissimilarity metric, $D$, calculations
@@
The proposed method is next demonstrated on a well-known mathematical function that is used in optimization studies. Ackley’s function is a customiziable multi-dimensional function that has been implemented in numerous optimization applications for testing and validation purposes [ackley2012connectionist, ackley2]. Here, 2D Ackley’s function, Equation 10, is implemented to further demonstrate the proposed interpretable multi-source data fusion method. In addition to the original 2D function, three (3) other variations of Ackley’s function are generated from the original function as different sources of information. Similar to the parabola problem, the original function is named as the Ground Source and the remaining 3 functions are named as Perturbed Sources 1, 2, and 3. The formulations of the functions are found in Supplementary Table 2 and their surface of the response, $z$, is shown in Figure 4a.
@@
Identical to the previous case study, two ML models are built (GP and LVGP) to demonstrate the effectiveness of the proposed approach. First, a GP model is built by only using the quantitative variables, $x$ and $y$ to predict the response, $z$. Then, in addition to the two quantitative variables, a qualitative variable is added to differentiate the information sources for the LVGP model. This approach is the exact same approach taken for the previous parabola problem and will be implemented for the remainder of this work. During the training, 5-fold cross validation (CV) technique is implemented and tested on a separate testing set. Both the training and testing sets are created by sampling uniformly from the space $[x,y]\epsilon[-5,5]$. The number of training and testing samples employed to build and validate the models are available in Supplementary Table 2. Next, a comparison study between the predictive performance of both models using the NMRSE metric is carried out. Table 2 (Model Types: GP and LVGP) shows the performance metrics for both models. From the results, it can be clearly observed once again that incorporating different sources as a separate variable significantly improves the predictive performance of the model. The model was also able to capture the relationships between the response surfaces and provide interpretability through the latent variables clearly. The latent space obtained from the LVGP, Figure 4b, matches very well with the response surfaces, Figure 4a. Furthermore, the dissimilarity metric calculated using the Ground Source as the reference source, $D$, agree with the response surfaces and what is known regarding the similarity between the functions, demonstrating the efficacy of the method and the metric (Figure 4b).
@@
A possible follow-up question one might ask as a response to the demonstrated results could be to build separate GP models to learn the information sources and predict their responses individually to achieve better accuracies. To answer this question, a separate GP model is built only using the Ground Source and compared with the previously built LVGP on a separate testing set. For ease of notation, this single-source model is denoted as GP-GS. The trained GP-GS model only contained 20 samples from the Ground Source whereas the LVGP model contained all 170 samples from the available sources. The same testing set from the Ground Source is used and the prediction performances are compared. Table 2 (Model Types: LVGP and GP-GS) shows the prediction performances between the GP-GS and LVGP models, along with the parity plot in Figure 4c. In addition to achieving lower NRMSE values for accurate modeling, the parity plot demonstrates that the prediction uncertainty provided by the LVGP model is much less compared to the GP model. This result reveals an interesting phenomenon. Compared to the model that is only trained on the original single-source of information, gathering further information from other sources and explicitly denoting information sources into the data fusion modeling results in a more confident predictive model that provides lower prediction uncertainty on the original single-source of information.
@@
Table 2: Predictive performance comparison between GP, LVGP, and GP-GS models on the 2D Ackley function
Model   Type Mean Training NRMSE Mean CV NRMSE Testing NRMSE
GP $0.253$ $0.289$ $0.254$
LVGP $0.013$ $0.029$ $0.024$
GP-GS $0.089$
@@
The conclusions from this study can be significant for many reasons and can be applied to many engineering and scientific applications, specifically for materials science. This is because many researchers only have access to or control over their own experiments/simulations and rely on building a surrogate model using their own data to guide their next property evaluations. On the other hand, due to the high costs of property evaluations, data curation for better predictive surrogate models could be expensive and limited to only a few samples. However, by incorporating more data from different available sources (open-sources, papers, patents) and explicitly differentiating each data source as an input to their models, researchers have the potential to obtain better and confident predictive model, which in return has the potential to lead to novel material candidate discovery at a faster rate. In the following two sections, Section 4 and Section 5, the advantages of using multi-source data fusion through LVGP on two materials design cases are demonstrated.
@@
Figure 4: (a) Response Surface of the Ground Source and Perturbed Sources for the 2D Ackley’s function. (b) The latent variables obtained from the LVGP model. Ground Source is denoted as the reference source (*) for dissimilarity metric, $D$, calculations. (c) The parity plot between the GP-GS and LVGP models tested on the Ground Source data only